OpenStudy (anonymous):

a^2+-9 over a^2 times a^2+-3a over a^2+a+-12

OpenStudy (anonymous):

\[\frac{a^2-9}{a^2}\times \frac{a^2-3a}{a^2+a-12}\] \[\frac{(a+3)(a-3)a(a-3)}{a^2(a+4)(a-3)}\] \[\frac{(a+3)(a-3)}{a(a+4)}\]